OpenStudy (anonymous):

1 3/7 *2/6 SOLVE N SHOW WORK

OpenStudy (anonymous):

10/7 * 2/6 20/42 10/21

OpenStudy (anonymous):

thank u could u do another so that i understand fully

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

7 6/9* 1/3

OpenStudy (anonymous):

69/9 * 1/3 23/9

OpenStudy (anonymous):

so what do i do when the problem look like 1 6/12* 7 2/46

OpenStudy (anonymous):

what do you think? what have i been doing? just do exactly the same thing

OpenStudy (anonymous):

change it to improper form and multiply and then simplify

OpenStudy (anonymous):

ok thanks but please dont get upset i m only n the 5th grade

OpenStudy (anonymous):

i'm not getting upset? oO i'm just saying think about it

OpenStudy (anonymous):

do you know how to get it into improper form?

OpenStudy (anonymous):

i * the bottom number with the whole number then i add the top number once i have done that i do the same with the other side then i * the 2 fractions 2gether right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do you want to try it?

OpenStudy (anonymous):

sure give me a problem

OpenStudy (anonymous):

how bout that one that you just gave me 1 6/12* 7 2/46

OpenStudy (anonymous):

ok i got 5832/552

OpenStudy (anonymous):

that's right, but you need to simplify it do you know how?

OpenStudy (anonymous):

trying 2 do that now

OpenStudy (anonymous):

i think i got it 243/23

OpenStudy (anonymous):

and also, those are huge numbers, you can make your life a lot easier if you simplify before you multiply, like this \[\frac{18}{12} * \frac{324}{46} = \frac{3}{2}*\frac{162}{23} = \frac{3}{1}*\frac{81}{23} = \frac{243}{23}\] you got it right if you simplify before you multiply, it will be easier to get the answer make sense?

OpenStudy (anonymous):

i got it thank u 4 all the help

OpenStudy (anonymous):

no problem good lucks